OpenStudy (anonymous):

y = x^2 − 2x − 3 Find the vertex. Find the x-intercepts. Describe the graph.

OpenStudy (anonymous):

Let's see...for the vertex "-b/2a" so that's 2/2(1) = 1 (Vertex)

OpenStudy (anonymous):

the vertex is at the point where the derivative or slope of this function is zero so dy/dx=2x-2 2x-2=0 x=1 so substituting in f(x) we get 1-2-3=-4 so vertex at (1,-4) as the function is in x and the coefficient of x is positive the graph is like 

OpenStudy (anonymous):

we get x intercept when the graph touches x axis so putting y=0 in the function we get x^2-2x-3=0 now solve the quadratic equation to get the x intercept